OpenStudy (studybird123):

A box contains 12 green balls, 11 red balls, 7 white balls, and 10 black balls. What is the probability of drawing a black ball from this box?

OpenStudy (studybird123):

A.1/40 B. 1/10 C. 1/4 D.1/3

OpenStudy (anonymous):

probability = (success options)/(all options) = (10 black) / (40 in all) = answer

OpenStudy (anonymous):

add all up and do 10 divided by the number reduced

OpenStudy (anonymous):

c

OpenStudy (anonymous):

10/40=1/4